# Stackelberg vs. Nash in the Lottery Colonel Blotto Game

### v1

@@ 4 The Leader’s Optimal Commitment Strategy
Theorem 1 characterizes $K(\bm{x}_{a})$, the support of the follower’s best response strategy when the leader employs their optimal commitment strategy $\bm{x}_{a}$. In this section, we formulate the leader’s objective as an optimization problem for any given set $K(\bm{x}_{a})$. By solving the optimization problem for all $n$ possible realizations of the set $K(\bm{x}_{a})$ as described in Theorem 1, and comparing the optimal values, we identify the optimal commitment strategy $\bm{x}_{a}$.
Assume the support of the follower’s best response strategy, when the leader employs their optimal commitment strategy, is given by $K$. Given a specific set $K(\bm{x}_{a})$, the leader’s optimal commitment strategy can be determined by solving the following optimization problem (OC).
[…]
@@
$x_{bj}=\frac{(x_{aj}\cdot v_{bj})^{\frac{1}{2}}\cdot(\sum_{j^{\prime}\in K}x_{aj^{\prime}}+x_{b})}{\sum_{j^{\prime}\in K}(x_{aj^{\prime}}\cdot v_{bj^{\prime}})^{\frac{1}{2}}}-x_{aj}>0,\forall j\in K,$ (9)
@@
The leader’s utility is derived from two components: the battlefield set $K$, where the leader and the follower share the battlefield proportional to their resources allocated, and the set $\overline{K}$, where the leader has completely won the battlefield. Equation (7) represents the leader’s budget constraint. Inequalities (8) assume the leader allocates a positive amount of budget to every battlefield in the optimal commitment strategy. Constraints (9) and (10) describe the follower’s best response strategy as outlined in Lemma 1.
@@
Lemma 1 provides us with $\frac{v_{bj}}{x_{aj}}\leq\frac{\left(\sum_{l\in K(\bm{x}_{a})}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}{\left(x_{b}+\sum_{l\in K(\bm{x}_{a})}x_{al}\right)^{2}}$, $\forall j\in\overline{K(\bm{x}_{a})}$. This implies $x_{aj}\geq\frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K(\bm{x}_{a})}x_{al}\right)^{2}}{\left(\sum_{l\in K(\bm{x}_{a})}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}}$. To demonstrate that equality holds for any $j\in\overline{K(\bm{x}_{a})}$, we compare the leader’s utility before and after shifting an arbitrarily small amount of budget $\epsilon$ from a battlefield $j\in\overline{K(\bm{x}_{a})}$ to a battlefield $j\in K(\bm{x}_{a})$. If this budget shift consistently increases the leader’s utility, the leader will allocate precisely the amount stated in the lemma to keep the set $\overline{K(\bm{x}_{a})}$ unchanged. The full proof of the lemma is deferred to the Appendix B.
Then, we simplify the optimization problem (OC) in three aspects: (i) we add the expression for $x_{aj},j\in\overline{K}$ as a constraint in (OC); (ii) we substitute $x_{bj}$ as outlined in (9) and (10) into the objective function (6), thereby removing $x_{bj}$ from the leader’s optimization problem; (iii) we note that the second term of (6) can be removed, as it becomes a constant when the set $\overline{K}$ is fixed. Denote the objective $\hat{u}_{a}(\bm{x}_{a})=u_{a}(\bm{x}_{a},\bm{x}_{b})-\sum_{j\in\overline{K}}v_{aj}$.
Therefore, given the set $K(\bm{x}_{a})$, the optimization problem (OC) can be reformulated as follows.
[…]
@@
For the sake of simplicity, we define the following expressions.
Definition 1.
For $i\in\{a,b\}$, $K\subseteq[n]$, let
$v_{iK}$ $=\sum_{j\in K}v_{ij},$
$c_{K}$ $=\sum_{j\in K}\frac{v_{aj}^{2}}{v_{bj}},$
$B_{1}(K)$ $=x_{a}v_{bK}^{2}-2x_{b}v_{bK}v_{b\overline{K}},$
$B_{2}(K)$ $=4x_{b}v_{aK}v_{b\overline{K}}-2x_{a}v_{aK}v_{bK},$
$B_{3}(K)$ $=x_{a}v_{aK}^{2}-2x_{b}v_{b\overline{K}}c_{K},$
$B_{4}(K)$ $=x_{a}^{2}v_{bK}^{2}-4x_{b}(x_{a}+x_{b})v_{bK}v_{b\overline{K}},$
$B_{5}(K)$ $=8x_{b}(x_{a}+x_{b})v_{aK}v_{b\overline{K}}-2x_{a}^{2}v_{aK}v_{bK},$
$B_{6}(K)$ $=x_{a}^{2}v_{aK}^{2}-4x_{b}(x_{a}+x_{b})c_{K}v_{b\overline{K}},$
$\phi_{1}(\theta,K)$ $=B_{1}(K)\theta^{2}+B_{2}(K)\theta+B_{3}(K),$
$\phi_{2}(\theta,K)$ $=B_{4}(K)\theta^{2}+B_{5}(K)\theta+B_{6}(K).$
@@
assume, without loss of generality, that the battlefields are ordered such that $\frac{v_{a1}}{v_{b1}}\leq\frac{v_{a2}}{v_{b2}}\leq\cdots\leq\frac{v_{an}}{v_{bn}}$. For each set $K\in\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$, solving the optimization problem (OC’) can yield the leader’s candidate optimal commitment strategy $\bm{x}_{a}$ as follows.
Case 1. $\forall j,h\in K$, $\frac{v_{aj}}{v_{bj}}=\frac{v_{ah}}{v_{bh}}$, or $K=\{1\}$, then
$x_{aj}=\begin{cases}x_{aK}\cdot\frac{v_{aj}}{v_{aK}},&\text{if}\;j\in K;\\ \frac{v_{bj}\cdot\left(x_{b}+x_{aK}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},&\text{if}\;j\in\overline{K},\end{cases}$
where $x_{aK}=\frac{x_{a}v_{bK}-2x_{b}v_{b\overline{K}}+\sqrt{x_{a}^{2}v_{bK}^{2}-4x_{a}x_{b}v_{bK}v_{b\overline{K}}-4x_{b}^{2}v_{bK}v_{b\overline{K}}}}{2(v_{b\overline{K}}+v_{bK})}$.
Case 2.1. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K=[n]$, then
$x_{aj}=\frac{\left(\frac{v_{aj}}{\sqrt{v_{bj}}}+(\frac{c_{K}}{v_{bK}})^{\frac{1}{2}}\cdot\sqrt{v_{bj}}\right)^{2}}{\sum_{l=1}^{n}\left(\frac{v_{al}}{\sqrt{v_{bl}}}+(\frac{c_{K}}{v_{bK}})^{\frac{1}{2}}\cdot\sqrt{v_{bl}}\right)^{2}}\cdot x_{a},\quad\forall j\in[n].$
Case 2.2. there exist $j,h\in K$, such that $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, and $K\subsetneq[n]$, then
$x_{aj}=\begin{cases}\frac{2x_{b}^{2}v_{b\overline{K}}\cdot(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\cdot\sqrt{\phi_{2}(\alpha,K)}},\quad&j\in K;\\ \frac{v_{bj}\cdot\left(x_{b}+\sum_{l\in K}x_{al}\right)^{2}}{\left(\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}\right)^{2}},\quad&j\in\overline{K},\end{cases}$
where $\alpha$ is the optimal solution to the following optimization problem
$\max_{\alpha}\quad$ $\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{v_{bj}}-\alpha)\cdot v_{aj}\right)\cdot\left(\sum_{j\in K}(v_{aj}-\alpha v_{bj})\right)}{\frac{\phi_{1}(\alpha,K)-(v_{aK}-v_{bK}\alpha)\cdot\sqrt{\phi_{2}(\alpha,K)}}{2x_{b}^{2}v_{b\overline{K}}}\cdot x_{b}+\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}}$
s.t. $x_{a}^{2}-4\cdot\frac{\left(\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}\right)\cdot v_{b\overline{K}}}{\left(\sum_{h\in K}(v_{ah}-\alpha v_{bh})\right)^{2}}\cdot x_{b}(x_{a}+x_{b})\geq 0.$
Proof.
Rearranging Equation (14), we obtain
Second, we consider the case where $\frac{v_{aj}}{v_{bj}}\neq\frac{v_{ah}}{v_{bh}}$, leading to subcases 2.1 and 2.2.
In this case, $\beta\neq 0$. According to Lemma 3, for $\forall j\in K$, we have $(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha v_{bj}^{\frac{1}{2}})\frac{1}{\beta}=x_{aj}^{\frac{1}{2}}$. The resources invested by the leader in $K$ are $\frac{1}{\beta^{2}}\sum_{j\in K}(\frac{v_{aj}}{\sqrt{v_{bj}}}-\alpha\sqrt{v_{bj}})^{2}$. For $\forall j\in\overline{K}$, we have
We will analyze different values of $K$ and leave this proof to the Appendix B. ∎
The Leader’s Optimal Commitment Strategy. In a Stackelberg game $\mathcal{G}:=\langle\{a,b\},[n],x_{a}$, $x_{b}$, $(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle$, following Theorem 2, we can compute the candidate optimal commitment strategy $\bm{x}_{a}$ for each set $K(\bm{x}_{a})\in\{\{1\},\{1,2\},\dots,\{1,2,\dots,n\}\}$. For each candidate strategy $\bm{x}_{a}$, we then determine the leader’s utility. The strategy $\bm{x}_{a}$ that yields the highest utility is identified as the leader’s optimal commitment strategy.
Recall that the best response function provided by Kovenock and Arjona (2019) requires $x_{aj}>0$, for all $j\in[n]$. To conclude this section, we demonstrate that if there exists a battlefield $j$ such that $x_{aj}=0$ in the leader’s optimal commitment strategy, as identified through our process, then increasing $x_{aj}$ to an arbitrarily small budget and carefully adjusting resource allocation in other battlefields will enhance the leader’s utility. Therefore, the leader’s optimal commitment strategy must satisfy this prerequisite.
@@
Lemma 5 allows us to merge battlefields with identical relative value ratios. This step is essential for proving the main result of this section, as detailed below. First, for simplicity, let’s define a few expressions.
Definition 2.
For any $T\subseteq[n]$, $i\in\{a,b\}$, let
$v_{iT}$ $=\sum_{j\in T}v_{ij},$
$t_{1}(T)$ $=\Psi_{1}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}}),$
$t_{2}(T)$ $=\Psi_{1}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT}),$
$t_{3}(T)$ $=\Psi_{2}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT},t_{2}(T),t_{1}(T))$
$\quad+\Psi_{3}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}},t_{1}(T),t_{2}(T)),$
$R(T)$ $=\frac{t_{4}(T)}{t_{3}(T)}.$
@@
A Stackelberg equilibrium is also a Nash equilibrium if and only if one of the following conditions hold:
@@
  • There exists a set $M$, such that
    (1) $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for any $j,k\in M$, and $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for any $j,k\in\overline{M}=[n]\backslash M$, and (2) $\frac{x_{a}}{x_{b}}=R(T)$.
Note that the leader can always employ the Nash Equilibrium strategy in the Stackelberg game to ensure that their payoff is at least equal to that in the normal form game. Theorem 3 characterizes the scenarios under which being the first mover does not provide any additional benefit.
6 Leader’s Advantage and Impact
In this section, we analyze the advantages of the leader as a first mover, with these advantages being contingent upon the leader’s budget relative to that of the follower. Intuitively, when the leader makes an optimal commitment, the leader’s utility increases, while the follower’s utility tends to decrease. However, we are surprised to find that the follower’s utility can still improve in response to the leader’s optimal commitment.
Player’s strategy in the Nash equilibrium. Li and Zheng (2022) present an approach to compute the Nash equilibrium in two-player Lottery Colonel Blotto game. We can utilize this approach to compute Nash equilibria.
Lemma 6.
The leader’s budget is relatively large. The following theorem gives a lower bound for the leader’s utility in Nash equilibria.
@@ 6 Leader’s Advantage and Impact
It immediately shows that the utility of the leader in the Nash equilibrium is greater than or equal to $\frac{x_{a}}{x_{a}+x_{b}}\cdot(\sum_{j=1}^{n}v_{aj})$. ∎
@@
Let $u_{a}^{NE}$ and $u_{a}^{SE}$ denote the leader’s utility under the Nash equilibrium and Stackelberg equilibrium, respectively. We have $\frac{u_{a}^{SE}}{u_{a}^{NE}}\leq\frac{x_{a}+x_{b}}{x_{a}}$
@@
This corollary shows that when $x_{a}$ is relatively large, $\frac{u_{a}(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})}{u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})}$ approaches 1. Hence, when the leader’s budget is relatively large, the leader can make a commitment to enhance their own benefits, but the improvement is marginal.
The ratio of the leader’s utility under the Stackelberg equilibrium to that under the Nash equilibrium. We consider a game with two battlefields. Analyzing this two-battlefield scenario can provide insights into the leader’s advantages in games with multiple battlefields. Let $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},x_{a},1,(v_{a1},1),(v_{b1},1)\rangle$. Without loss of generality, let $v_{a1}\leq v_{b1}$. The following theorem gives the range of the ratio.
@@
In game $\tilde{\mathcal{G}}$, let $\hat{\bm{x}}_{a}$ denote the leader’s optimal commitment, $\hat{\bm{x}}_{b}$ denote the follower’s best response, and $u_{a}^{NE}$ denote the leader’s utility in the Nash equilibrium. Then, the ratio of Player $a$’s utility in the Stackelberg equilibrium to their utility in the Nash equilibrium is bounded by
@@
The leader’s budget is relatively small. Consider an extreme case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},\epsilon,1,(\epsilon,1),(o(\epsilon),1)\rangle$. We observe that when the leader’s budget is dominated by the follower’s budget, and the follower’s value on the first battlefield is negligible compared to the leader’s value on the first battlefield, the ratio of the leader’s utility when moving first to their utility in the Nash equilibrium approaches infinity. This suggests that the player with weak budget has a stronger incentive to act as the first mover.
The changes in the follower’s utility. When the leader makes an optimal commitment, it is natural to think that the follower’s utility would decrease. However, we find that this is not always the case. The follower’s utility may either decrease or increase.
Consider the case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},40,1,(0.01,1),(0.1,1)\rangle$. The case has unique Nash equilibrium, where $\bm{x}_{a}^{*}=(2.884,37.116)$ and $\bm{x}_{b}^{*}=(0.437,0.563)$, resulting in player $a$’s utility being 0.994 and player $b$’s utility being 0.028. In the Stackelberg equilibrium, $\bm{x}_{a}=(1.128,38.872)$ and $\bm{x}_{b}=(0.968,0.032)$, with the leader’s utility rising to 1.005 and the follower’s utility increasing to 0.047. The respective ratios are $\frac{1.005}{0.994}=1.011$ for the leader and $\frac{0.047}{0.028}=1.679$ for the follower.
It is somewhat unexpected that both players’ utilities in the Stackelberg equilibrium exceed their payoffs in the Nash equilibrium. However, this outcome is entirely plausible, as this is a non-zero-sum game. Since the leader values the second battlefield more, they allocate more of their budget to it, minimizing the follower’s share. Meanwhile, the follower values the first battlefield more, and the leader’s strategy incidentally increases the follower’s share of the first battlefield, improving the follower’s overall payoff.
Consider another case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},0.1,1,(2,1),(10,1)\rangle$. The unique Nash equilibrium is $\bm{x}_{a}^{*}=(0.036,0.064)$ and $\bm{x}_{b}^{*}=(0.738,0.262)$, where player $a$’s utility is 0.289 and player $b$’s utility is 10.338. In the Stackelberg equilibrium, the leader’s allocation is $\bm{x}_{a}=(0.063,0.037)$ and the follower’s is $\bm{x}_{b}=(0.822,0.178)$, yielding a leader utility of 0.315 and a follower utility of 10.118. The respective ratios are $\frac{0.315}{0.289}=1.090$ for the leader and $\frac{10.118}{10.338}=0.979$ for the follower.
In this case, since both players have similar preferences for the battlefields, the leader allocates more budget to the first battlefield to reduce the follower’s share and increase their own payoff. As a result, the follower’s share in the first battlefield decreases, leading to a reduction in their payoff from that battlefield, which in turn causes a decline in their overall payoff.

### v3

@@ 4 The Leader’s Optimal Commitment Strategy
Theorem 1 characterizes $K(\bm{x}_{a})$, the support of the follower’s best response strategy when the leader employs its optimal commitment strategy $\bm{x}_{a}$. In this section, we formulate the leader’s objective as an optimization problem for any given set $K(\bm{x}_{a})$. By solving the optimization problem for all $n$ possible realizations of the set $K(\bm{x}_{a})$ as described in Theorem 1, and comparing the optimal values, we identify the optimal commitment strategy $\bm{x}_{a}$.
Assume the support of the follower’s best response strategy, when the leader employs its optimal commitment strategy, is given by $K$. Given a specific set $K(\bm{x}_{a})=K$, i.e., the set of battlefields in which the follower participates when playing a best response, the leader’s optimal commitment strategy can be determined by solving the following optimization problem (OC).
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\quad$ $u_{a}(\bm{x}_{a},\bm{x}_{b})=\sum_{j\in K}\frac{x_{aj}\cdot v_{aj}}{x_{aj}+x_{bj}}+\sum_{j\in\overline{K}}v_{aj}$ (6)
@@
$x_{bj}=\frac{(x_{aj}v_{bj})^{\frac{1}{2}}(x_{b}+\sum\limits_{j^{\prime}\in K}x_{aj^{\prime}})}{\sum_{j^{\prime}\in K}(x_{aj^{\prime}}\cdot v_{bj^{\prime}})^{\frac{1}{2}}}-x_{aj}>0,\forall j\in K,$ (9)
$(\frac{v_{bj}}{x_{aj}})^{\frac{1}{2}}\leq\frac{\sum_{l\in K}(x_{al}v_{bl})^{\frac{1}{2}}}{x_{b}+\sum_{l\in K}x_{al}},\forall j\notin K,$ (10)
@@
The leader’s utility is derived from two components: the battlefield set $K$, where the leader and the follower share the battlefield proportional to their resources allocated, and the set $\overline{K}$, where the leader has completely won the battlefield. Equation (7) represents the leader’s budget constraint. Inequalities (8) assume the leader allocates a positive amount of budget to every battlefield in the optimal commitment strategy. Constraints (9) to (11) describe the follower’s best response strategy as outlined in Lemma 1.
@@
We simplify the optimization problem (OC) in three aspects: (i) we add the expression for $x_{aj},j\in\overline{K}$ as a constraint in (OC); (ii) we substitute $x_{bj}$ as outlined in (9) and (11) into the objective function (6), thereby removing $x_{bj}$ from the leader’s optimization problem; (iii) we note that the second term of (6) can be removed, as it becomes a constant when the set $\overline{K}$ is fixed. Denote the objective $\hat{u}_{a}(\bm{x}_{a})=u_{a}(\bm{x}_{a},\bm{x}_{b})-\sum_{j\in\overline{K}}v_{aj}$.
Therefore, given the set $K(\bm{x}_{a})=K$, the optimization problem (OC) can be reformulated as (OC’), as shown below.
$\max_{\bm{x}_{a}\in\bm{X}_{a}}\quad$ $\hat{u}_{a}(\bm{x}_{a})=\left(\sum_{j\in K}(\frac{x_{aj}}{v_{bj}})^{\frac{1}{2}}v_{aj}\right)\left(\frac{\displaystyle\sum_{j\in K}(x_{aj}v_{bj})^{\frac{1}{2}}}{x_{b}+\sum\limits_{j\in K}x_{aj}}\right)$ (12)
@@
for each set $K\in\{\{1\},\{1,2\},\cdots,\{1,2,\cdots,n\}\}$, we can, in $O(n)$ steps, formulate an optimization problem involving a univariate continuous function defined over the union of two half-open intervals. The solution to this problem represents the leader’s optimal commitment strategy.
Due to the complexity of the expression of the leader’s optimal commitment, we provide its specific formula in the proof of Theorem 2. Since the optimization objective for each $K$ is a univariate continuous function, this allows us to perform simulation experiments.
The Leader’s Optimal Commitment Strategy. Following Theorem 2, in a Stackelberg game $\mathcal{G}:=\langle\{a,b\},[n],x_{a},$ $x_{b},(v_{aj})_{j=1}^{n},(v_{bj})_{j=1}^{n}\rangle$, we compute the candidate optimal commitment strategy $\bm{x}_{a}$ for each set $,K(\bm{x}_{a})\in\{\{1\},\{1,2\},$$\dots,\{1,2,\dots,n\}\}$. For each candidate strategy $\bm{x}_{a}$, we then determine the leader’s utility. The strategy $\bm{x}_{a}$ that yields the highest utility is identified as the leader’s optimal commitment strategy.
Recall that the best response function provided by (Kovenock and Arjona, 2019) requires $x_{aj}>0$, for all $j\in$ $[n]$. To conclude this section, we demonstrate that if there exists a battlefield $j$ such that $x_{aj}=0$ in the leader’s optimal commitment strategy, as identified through our process, then increasing $x_{aj}$ to an arbitrarily small budget and carefully adjusting resource allocation in other battlefields will enhance the leader’s utility. Therefore, the leader’s optimal commitment strategy must satisfy this prerequisite.
@@
Lemma 5 allows us to merge battlefields with identical relative value ratios. This step is essential for proving the main result of this section, as detailed below.
Figure 1: Utility Curves in Two Equilibria.
@@
There exists a function $f:\mathbb{R}^{4}\rightarrow\mathbb{R}$ such that a Stackelberg equilibrium is also a Nash equilibrium if and only if one of the following conditions hold:
@@
  • There exists a set $M\subsetneq[n]$ and its complement $\overline{M}$ such that (1) $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for $\forall j,k\in M$, and $\frac{v_{aj}}{v_{bj}}=\frac{v_{ak}}{v_{bk}}$ for $\forall j,k\in\overline{M}$, and (2) $\frac{x_{a}}{x_{b}}=f\left(\sum_{j\in M}v_{aj},\sum_{j\in M}v_{bj},\sum_{j\in\overline{M}}v_{aj},\sum_{j\in\overline{M}}v_{bj}\right)$.
The function $f$ in Theorem 3 is complicated, so we provide its specific expression in the proof. To reinforce the understanding of Theorem 3, particularly the second case, we construct a class of instances and demonstrate how the utilities of both players vary across the two equilibria as their relative budgets change, as shown in Figure 1.
Let the budgets of the two players be $r$ and $1$, respectively, with their relative budget ratio defined as $r=\frac{x_{a}}{x_{b}}$. Let $v_{a1}=1$, $v_{a2}=5$, $v_{b1}=1$, and $v_{b2}=0.5$. Clearly, this game satisfies the second case of Theorem 3. In Figure 1, $SEu_{a}$ and $SEu_{b}$ represent the utilities of players $a$ and $b$ in the Stackelberg equilibrium, while $NEu_{a}$ and $NEu_{b}$ denote their utilities in the Nash equilibrium. As $r$ changes, the utilities of the two players in both equilibria also change. Notably, there exists a value of $r$ such that $r=\frac{x_{a}}{x_{b}}=f(1,1,5,0.5)$, at which both players achieve the same utility in both equilibria.
Furthermore, several interesting observations emerge from these instances. When $\frac{x_{a}}{x_{b}}>f$, both players experience higher utilities in the Stackelberg equilibrium than in the Nash equilibrium. In other words, an increase in the leader’s utility does not necessarily reduce the follower’s utility. Specifically, the leader’s optimal commitment in the Stackelberg equilibrium allocates more resources to the higher-value battlefield (Battlefield 2) and fewer resources to the lower-value battlefield (Battlefield 1) compared to the Nash equilibrium. Since the follower values the battlefields in the opposite way, they prefer Battlefield 1. This allows the follower to derive more utility from Battlefield 1, leading to an increase in their utility under the Stackelberg equilibrium. Conversely, when the leader’s budget is relatively small (i.e., when the ratio is less than $f$), the situation reverses. This is illustrated by the following calculation:
  • If $\frac{x_{a}}{x_{b}}=0.5$, which is less than $f$, the unique Nash equilibrium is $\bm{x}_{a}^{*}=(0.025,0.475)$, $\bm{x}_{b}^{*}=(0.340,0.660)$, where $NEu_{a}=2.161$, $NEu_{b}=1.223$. The Stackelberg equilibrium is $\bm{x}_{a}=(0.136,0.364)$, $\bm{x}_{b}=(0.559,0.441)$, where $SEu_{a}=2.458$, $SEu_{b}=1.079$.
  • If $\frac{x_{a}}{x_{b}}=2$, which is greater than $f$, the unique Nash equilibrium is $\bm{x}_{a}^{*}=(0.667,1.333)$, $\bm{x}_{b}^{*}=(0.833,0.167)$, where $NEu_{a}=4.889$, $NEu_{b}=0.611$. The Stackelberg equilibrium is $\bm{x}_{a}=(0.543,1.457)$, $\bm{x}_{b}=(0.847,0.153)$, where $SEu_{a}=4.915$, $SEu_{b}=0.657$.
6 Leader’s Advantage
In this section, we analyze the advantages of the leader as a first mover, with these advantages being contingent upon the leader’s budget relative to that of the follower. Specifically, we consider three cases: (1) the number of battlefields $n>2$, we find that when the leader’s budget is relatively large, the advantages of the leader is marginal; (2) $n=2$, we provide an upper bound and a lower bound on the ratio of the leader’s utility under the Stackelberg equilibrium to that under the Nash equilibrium; and (3) $n=2$ and the leader’s budget is relatively small, we find that the advantages of the leader could be infinite.
There are $n>2$ battlefields. The following theorem gives a lower bound for the leader’s utility in Nash equilibria.
@@ 6 Leader’s Advantage
It shows that the utility of the leader in the Nash equilibrium is greater than or equal to $\frac{x_{a}}{x_{a}+x_{b}}\cdot(\sum_{j=1}^{n}v_{aj})$. ∎
By Theorem 4, we have the following corollary.
@@
Let $u_{a}^{NE}$ and $u_{a}^{SE}$ denote the leader’s utility under the Nash equilibrium and Stackelberg equilibrium, respectively. We have $\frac{u_{a}^{SE}}{u_{a}^{NE}}\leq\frac{x_{a}+x_{b}}{x_{a}}$.
@@
This corollary shows that when $x_{a}$ is relatively large, $\frac{u_{a}(\tilde{\bm{x}}_{a},\tilde{\bm{x}}_{b})}{u_{a}(\bm{x}^{*}_{a},\bm{x}^{*}_{b})}$ approaches 1. Hence, when the leader’s budget is relatively large, the leader can make a commitment to enhance its own benefits, but the improvement is marginal.
There are two battlefields. We consider a game with two battlefields. Analyzing two-battlefield scenario can provide insights into the leader’s advantages in games with multiple battlefields. Let $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},x_{a},1,(v_{a1},1),(v_{b1},$ 1)1$\rangle$1 ) ⟩. Without loss of generality, let $v_{a1}\leq v_{b1}$. For the game $\tilde{\mathcal{G}}$, Li and Zheng (2022) present an approach to compute Nash equilibria. We can utilize this approach to compute Nash equilibria. The following theorem gives the range of the ratio.
@@
In game $\tilde{\mathcal{G}}$, let $\hat{\bm{x}}_{a}$ denote the leader’s optimal commitment, $\hat{\bm{x}}_{b}$ denote the follower’s best response, and $u_{a}^{NE}$ denote the leader’s utility in the Nash equilibrium. Then, the ratio of Player $a$’s utility in the Stackelberg equilibrium to its utility in the Nash equilibrium is bounded by
@@
There are two battlefields and the leader’s budget is relatively small. By Theorem 5, we can derive the following corollary, which describes the leader’s advantage when the leader’s budget is relatively small.
Corollary 2.
In game $\tilde{\mathcal{G}}$, the ratio of the leader’s utility when moving first to its utility in the Nash equilibrium approaches infinity.
To illustrate this corollary, we consider an extreme case $\tilde{\mathcal{G}}:=\langle\{a,b\},\{1,2\},\epsilon,1,(\epsilon,1),(o(\epsilon),1)\rangle$. We observe when the leader’s budget is dominated by the follower’s budget, and the follower’s value on the first battlefield is negligible compared to the leader’s value on the first battlefield, the ratio of the leader’s utility when moving first to its utility in the Nash equilibrium approaches infinity.
Although this corollary focuses on two battlefields, using the reduction method presented in Section 3, we can extend the analysis to multiple battlefields. Therefore, this corollary holds for scenarios with more than two battlefields as well.